# PolicyPad: Collaborative Prototyping of LLM Policies

### v1

@@ 5.2.3. Interactive in-editor scenario widgets
Figure 5. Scenarios can be brought into the editor inline with the policy as interactive widgets via referencing the scenario’s title with the ’@’ symbol. Once in the editor, all users can click on it, view it in their scenario sidebar, and flag responses for group discussion.:
@@ 5.2.4. Drafting policy statements
Once a group reviews the heuristics at the top of the editor to ensure they have a common understanding of policy desiderata, they are ready to start drafting policy statements. These policies address oddities, concerns, and other noteworthy aspects of model behavior they observed and flagged in the scenarios.
@@ 5.2.5. Testing the policy with scenarios
Figure 6. A drafting block (directly above “Conversational tone”) can be added into the editor to draft experimental policies without affecting model behavior.:
@@
Figure 7. PolicyPad offers two ways to test the behavior of the policy-informed model against a scenario: (1) regenerating the latest AI message, or (2) continuing the conversation.:
@@
Figure 8. Upon saving the policy via the Snapshot policy button, PolicyPad (1) adds the policy to the version history and generates a title summarizing key changes, (2) conducts an automated heuristic evaluation of the policy, and (3) updates the latest responses to all scenarios.:
@@ 5.2.6. Scenario spotlights
PolicyPad then automatically analyzes the group’s edits in the context of the current policy and heuristics. It uses a reasoning LLM to suggest a policy statement designed to steer the model towards producing a response more similar to the edited version (Fig. 9 4). If the user accepts the suggestion, it will be integrated into the policy. This alternative way of indirectly editing a policy through editing the model response is inspired by prior work on synthesizing principles from edits (Petridis et al., 2024; Louie et al., 2024).
@@
Figure 9. Workflow for spotlight scenarios. (1) A user can spotlight an interactive scenario widget to expand it into a card that everyone in the editor can view. 2) The model’s response can be edited collaboratively. After saving the edited response (3) and shrinking the spotlight scenario back into an interactive widget (4), PolicyPad automatically analyzes the edits in the context of the existing policy and heuristics to suggest a new policy statement.:
@@ 5.2.7. Heuristics editor and evaluator
Every time the policy is saved, PolicyPad runs an automated heuristic evaluation to highlight any unsatisfied heuristics (Fig. 8 2). This automated evaluation is meant to draw attention to the heuristics and encourage discussion around them, rather than conclusively determining their fulfillment. Group members can easily override the automated decision if they agree on a different assessment.
Users continue to engage in this iterative process of policy drafting and experimentation until the session concludes.
@@ 5.3. Technical Details
PolicyPad is implemented as a web application built with React, TypeScript, and TipTap [^4] https://tiptap.dev/.. The real-time collaboration engine is supported via TipTap Cloud. PolicyPad uses serverless functions to call the OpenAI and Together.ai APIs. The policy-informed model is an instance of Llama 3.3 70B Instruct Turbo [^5] The ideal model for policy prototyping generates comparable responses to popular chatbot products (e.g., ChatGPT, Claude) but does not have an existing policy trained into it outside of basic guardrails. Llama 3.3 is an apt fit given its capable performance and bring-your-own safeguards setup (which we did not use). We also used Llama 3.3 instead of Llama 4 due to reports of the latter being narrowly optimized for specific benchmarks (Robison, 2025). hosted on Together.ai. The policy was fed into the model as a system prompt with some additional scaffolding to ensure the model followed it. We called GPT-4o for miscellaneous features that required light LLM processing (e.g., generating titles of new policy versions that capture key changes), and o4-mini for features that benefited from deeper reasoning (i.e. suggesting policy statements after response edits and automated heuristic evaluation).
@@ 6. Evaluation Study
To evaluate PolicyPad, we ran a series of group-based, within-subjects studies with 22 domain experts from two domains (10 from mental health, 12 from law). Our goal was to determine how the design decisions made for PolicyPad enhanced the policy prototyping experience. We also evaluate the outputs of PolicyPad by analyzing the policies created by experts to determine their novelty with respect to established LLM policies like Claude’s Constitution (Anthropic, 2023) and OpenAI’s Model Spec (OpenAI, 2025a) (as of September 1, 2025). Thus, we asked the following research questions:
@@
We selected mental health and law as our domains because they are regulated, high-stakes domains for which AI use has been increasing but contested [^6] OpenAI CEO Sam Altman observed that the younger generation uses it as a “therapist, a life coach […] asking, ‘What should I do?”’ despite the absence of confidentiality protections that govern human attorneys and therapists (Perez, 2025). (Lamparth et al., 2025; Moore et al., 2025; Cheong et al., 2024; Lohr, 2023). Crafting responsible LLM policies is therefore critical for ensuring users’ safety and well-being. The two domains are also distinct enough for us to observe how approaches to policy prototyping and the resulting policies can differ across domains.
@@ 6.1. Participants and Setup
We organized experts into small groups of 2–4 (median = 3). We observed during our formative study that groups of around three experts struck an ideal balance between creating a collaborative atmosphere and allowing room for meaningful individual contributions. Full participant and group details are available in Appendix G. Half the groups (4 of 8) contained participants who already knew each other from professional contexts. We did not observe this to impact the quality of discussions nor policies prototyped.
We facilitated policy prototyping sessions by providing (with order counterbalanced) PolicyPad and a baseline system to each group, followed by a brief exit interview. The total length of the study was 90 minutes. Participants were compensated \$150 USD in their choice of cash or a gift card upon completing the study. All studies were recorded and transcribed. This study was classified as exempt by our institution’s IRB.
@@ 6.1.2. Baseline system
Our baseline system implemented a simplified policy prototyping workflow that consisted only of iterative policy drafting and experimentation with a policy-informed model. It used the same collaborative policy editor as PolicyPad, but did not include built-in support for interactive scenarios (i.e, the scenario sidebar, scenario widgets, spotlight scenarios) nor heuristic evaluation. It resembled a more polished version of the first prototype used in Section 5.1: an editor with a policy-informed model in the sidebar (Fig. 11).
@@ 6.1.3. Starting materials
We prepared scenarios, heuristics, and a small amount of starter text for the policy. For each domain, the first author crafted 10 scenarios that represent a realistic conversation between a human user and an AI chatbot, using the same Llama 3.3 instance as PolicyPad to generate the responses. To maximize realism without access to product interaction logs, we sourced topics and language from datasets containing realistic questions in our domains of interest: MENTAT from Lamparth et al. (Lamparth et al., 2025) for mental health, and r/legaladvice-style cases from Cheong et al. (Cheong et al., 2024) for law. We varied the length of scenarios to be 1–5 conversational turns. For each group, we randomly sampled half the scenarios (5) to assign to the first task, and the rest were assigned to the second. In the system condition, scenarios were loaded directly into the system. In the baseline condition, the first user messages across the 5 scenarios were copied into a Google Doc and shared with participants upon request [^8] In the baseline condition, we gave participants the option of starting with or without first browsing these user messages..
Besides scenarios, we provided three basic heuristics to encourage clear and precise policy writing that draws from real-world professional practices. We also provided an Objectives section in the policy as examples of policy statements, drawn from objectives in OpenAI’s Model Spec (OpenAI, 2025a). The full starter heuristics and policy are available in Appendix E.
@@ 6.2. Procedure
  • Post-study survey: Group members swapped policies with another group in their domain and rated the policies on 5-point Likert scale questions (see Appendix F).
@@ 6.3.2. Policy novelty analysis (RQ2)
To analyze the novelty of policies—whether they contribute new perspectives, ideas, considerations, dependencies, or approaches to existing policies—we gathered experts’ policy statements from all sessions and evaluated each against publicly available policies for guiding responsible model behavior (the “existing set”). We combined all policies from OpenAI’s Model Spec (OpenAI, 2025a), Claude’s Constitution (Anthropic, 2023), and principles derived from workshops with legal experts in prior work (Cheong et al., 2024), to represent the set of existing policies. [^9] Excluding policies specifically targeted at moderating hate speech and disturbing content (e.g., (Lam et al., 2024)), as they are orthogonal to the policies we focus on in our sessions.
@@
    For each expert-written policy statement, we used GPT-4.1 with 3 prompts with varying definitions of novelty to make binary novelty decisions against the existing policies. Our prompts also required the LLM to generate a justification for its decision.
@@
    Any policies that were not unanimously determined to be novel in all 3 prompt evaluations were considered not novel. For the remaining policies, we further prompted GPT-4.1 to retrieve relevant quotes from the existing policies to be used as context for human evaluation.
@@
    Finally, two human annotators (members of our research team) reviewed the list of policies and quotes to make a final novelty determination. Annotations were first done independently, with any disagreements then resolved via a round of discussion. If annotators failed to reach a consensus, the policy statement was considered not novel by default.
While the reliability of LLMs for making content judgments has been called into question by recent work (Thakur et al., 2025; Chehbouni et al., 2025; Szymanski et al., 2025; Kim et al., 2025; Krumdick et al., 2025; Chehbouni et al., 2025), we structured our evaluation process to minimize the potential impacts of these factors. Specifically, we attempt to control for sensitivity to prompts by using multiple variations of prompts for initial evaluation, we request justifications and quote extraction to allow identification of possible model hallucinations, and we ensured the final novelty determination is done by human annotators. Overall, we believe that this process should yield a conservative determination of novelty, while also ameliorating challenges around human attention during review and comparison of exceptionally long texts.
@@ 7.1.1. Heuristics built common ground and inspired richer policies (system only)
Besides serving as guidelines, heuristics also served as entry points for deeper discussion on key policy topics. The starter heuristic on incorporating real-world professional practices into the policy initiated discussions in MH01, MH02, and MH04 about motivational interviewing (MI), a foundational technique for therapists. Experts then incorporated various aspects of MI into the policy, such as encouraging “summaries of conversation when appropriate” (MH01), and “Repeating or paraphrasing what [the user] is saying” (MH02). MH02 and MH03 brought up limits of confidentiality—when a mental health expert is legally or ethically required to break confidentiality share client information, even though expert-client conversations are otherwise private. MH02 brainstormed situations in which experts needed to break confidentiality (e.g., “immediate risk of harm to oneself or others; suicidal thoughts, urges, or behaviors; presence or risk of non-suicidal self injury”) and added a policy to “avoid using MI” in those situations. MH03 agreed that models, just like when experts work with clients, should provide a disclaimer early on in the conversation of conditions under which confidentiality will be broken and remind the user that “[the conversation] is not a confidential setting” when those conditions are triggered.
Interestingly, a couple groups of legal experts used the heuristic to debate whether real-world practices for lawyers and other legal professionals should even apply to AI, since they clearly established that AI does not have the same legal status as human lawyers. P17 in L02 shared that while “we lawyers do have rules of professional responsibility that we need to adhere to, they don’t apply to non-lawyers” and suggested removing that heuristic. P20 in L03 echoed that sentiment: “Those ethical rules of lawyers do not apply to AI systems.” Their group member, P19, agreed, and noted that “if ethical rules of lawyers did apply, then the AI model cannot even begin to suggest answers.” In general, this was an important point of distinction between the mental health and legal policies.
@@
Participants also expressed appreciation of the ability to edit the response and receive a system-generated policy suggestion. 100% of the policy suggestions were accepted during our studies. P17 shared that the suggestions were helpful in articulating their thoughts: “sometimes it’s difficult to put your thoughts into words, and the [suggestions] are helping you with that.” P10 agreed, saying that “the ability to pull that response in, edit it, and have a generated guardrail could be a huge time saver.” They saw as a way of removing the need for low-level wordsmithing: “We don’t need to edit that response perfectly, but if we can make it clear what our priorities are, and then see if the AI gets our nuance, that’s pretty incredible.” We also observed that in MH04, P9 and P10’s policy drafting began slowly, but accelerated considerably when they received policy suggestions that inspired more ideas. P7 shared why they accepted policy suggestions even when they seemed imperfect: “I liked the policy [suggestions], even if they weren’t necessarily dead on. It gave us ideas for other [policy statements].”

### v3

@@ 5.2.3. Interactive in-editor scenario widgets
Figure 6. Scenarios can be brought into the editor inline with the policy as interactive widgets via referencing the scenario’s title with the ’@’ symbol. Once in the editor, all users can click on it, view it in their scenario sidebar, and flag responses for group discussion.
Scenarios can be brought into the editor inline with the policy as interactive widgets via referencing the scenario’s title with the ’@’ symbol. Once selected, it appears inline in the editor alongside text and other widgets. Once in the editor, all users can click on it, view it in their scenario sidebar, and flag responses for group discussion. If flagged, the widget will glow orange for everyone.
@@ 5.2.4. Drafting policy statements
Once a group reviews the heuristics at the top of the editor to ensure they have a common understanding of desired policy goals, they are ready to start drafting policy statements. These policies address oddities, concerns, and other noteworthy aspects of model behavior they observed and flagged in the scenarios.
@@ 5.2.5. Testing the policy with scenarios
Figure 7. A drafting block (directly above “Conversational tone”) can be added into the editor to draft experimental policies without affecting model behavior.
A drafting block that is in a slightly lighter shade of gray than the rest of the text can be added into the editor to draft experimental policies without affecting model behavior.
@@
Figure 8. PolicyPad offers two ways to test the behavior of the policy-informed model against a scenario: (1) regenerating the latest AI message, or (2) continuing the conversation.
UI showing two ways to test the behavior of the policy-informed model against a scenario: (1) regenerating the latest AI message, or (2) continuing the conversation by sending any message.
@@
Figure 9. Upon saving the policy via the Snapshot policy button, PolicyPad (1) adds the policy to the version history and generates a title summarizing key changes, (2) conducts an automated heuristic evaluation of the policy, and (3) updates the latest responses to all scenarios.
Upon saving the policy via the Snapshot policy button, PolicyPad (1) adds the policy to the version history and generates a title summarizing key changes, (2) conducts an automated heuristic evaluation of the policy and checks off heuristics that are satisfied, and (3) updates the latest responses to all scenarios.
@@ 5.2.6. Scenario spotlights
PolicyPad then automatically analyzes the group’s edits in the context of the current policy and heuristics. It uses a reasoning LLM (o4-mini) to suggest a policy statement designed to steer the model towards producing a response more similar to the edited version (Fig. 10 4). If the user accepts the suggestion, it will be integrated into the policy. This alternative way of indirectly editing a policy through editing the model response is inspired by prior work on synthesizing principles from edits (Petridis et al., 2024; Louie et al., 2024).
@@
Figure 10. Workflow for spotlight scenarios. (1) A user can spotlight an interactive scenario widget to expand it into a card that everyone in the editor can view. 2) The model’s response can be edited collaboratively. After saving the edited response (3) and shrinking the spotlight scenario back into an interactive widget (4), PolicyPad automatically analyzes the edits in the context of the existing policy and heuristics to suggest a new policy statement.
UI workflow for spotlight scenarios. (1) A user can spotlight an interactive scenario widget to expand it into a card that everyone in the editor can view. (2) The model’s response can be edited collaboratively. After saving the edited response (3) and shrinking the spotlight scenario back into an interactive widget (4), PolicyPad automatically analyzes the edits in the context of the existing policy and heuristics to suggest a new policy statement, which is inserted into the editor. The user may choose to accept or reject it.
@@ 5.2.7. Heuristics editor and evaluator
Every time the policy is saved, PolicyPad runs an automated heuristic evaluation using o4-mini to highlight any unsatisfied heuristics (Fig. 9 2). This automated evaluation is meant to draw attention to the heuristics and encourage discussion around them, rather than conclusively determining their fulfillment. Group members can easily override the automated decision if they agree on a different assessment.
Users continue to engage in this iterative process of policy drafting and experimentation until the policy prototyping session concludes.
@@ 5.3. Technical Details
PolicyPad is implemented as a web application built with React, TypeScript, and TipTap. [^8] https://tiptap.dev/. The real-time collaboration engine is supported via TipTap Cloud. PolicyPad uses serverless functions to call the OpenAI and Together.ai APIs. The policy-informed model is an instance of Llama 3.3 70B Instruct Turbo [^9] The ideal model for policy prototyping generates comparable responses to popular chatbot products (e.g., ChatGPT, Claude) but does not have an existing policy trained into it outside of basic guardrails. Llama 3.3 is an apt fit given its capable performance and bring-your-own safeguards setup (which we did not use). We also used Llama 3.3 instead of Llama 4 due to reports of the latter being narrowly optimized for specific benchmarks (Robison, 2025). hosted on Together.ai. The policy was fed into the model as a system prompt with some additional scaffolding to ensure the model followed it. We called GPT-4o for miscellaneous features that required light LLM processing (e.g., generating titles of new policy versions that capture key changes), and o4-mini for features that benefited from deeper reasoning (i.e. suggesting policy statements after response edits and automated heuristic evaluation). Our prompts are available in our Supplementary Materials.
@@ 6. Evaluation Study
To evaluate PolicyPad, we ran a series of group-based, within-subjects studies with 22 domain experts from two domains (10 from mental health, 12 from law). Our goal was to determine how the design decisions made for PolicyPad enhanced the policy prototyping experience. We also evaluate the outputs of PolicyPad by analyzing the policies created by experts to determine their novelty with respect to established LLM policies like Claude’s Constitution (Anthropic, 2023) and OpenAI’s Model Spec (OpenAI, 2025b). Thus, we asked the following research questions:
@@
We selected mental health and law as our domains because they are regulated, high-stakes domains for which AI use has been increasing but contested [^10] OpenAI CEO Sam Altman observed that the younger generation uses it as a “therapist, a life coach […] asking, ‘What should I do?”’ despite the absence of confidentiality protections that govern human attorneys and therapists (Perez, 2025). (Lamparth et al., 2025; Moore et al., 2025; Cheong et al., 2024; Lohr, 2023). Crafting responsible LLM policies is therefore critical for ensuring users’ safety and well-being. The two domains are also distinct enough for us to observe how approaches to policy prototyping and the resulting policies can differ across domains. For each domain, we randomly sampled scenarios from datasets compiled by prior work in mental health (Lamparth et al., 2025) and law (Cheong et al., 2024). More details about scenario construction can be found in Section 6.1.3.
@@ 6.1. Participants and Setup
We organized experts into small groups of 2–4 (median = 3). We observed during our formative study that groups of around three experts struck an ideal balance between creating a collaborative atmosphere and allowing room for meaningful individual contributions. Full participant and group details are available in Appendix F. Half the groups (4 of 8) contained participants who already knew each other from professional contexts. We did not observe this to impact the quality of discussions nor policies prototyped.
We facilitated policy prototyping sessions by providing (with order counterbalanced) PolicyPad and a baseline system to each group, followed by a brief exit interview. The total length of the study was 90 minutes. Participants were compensated \$150 USD in their choice of cash or a gift card upon completing the study. All studies were recorded and transcribed. This study was classified as exempt by the University of Washington IRB.
@@ 6.1.2. Baseline system
Our baseline system implemented a simplified policy prototyping workflow that consisted only of iterative policy drafting and experimentation with a policy-informed model identical to the one used in PolicyPad. It used the same collaborative policy editor as PolicyPad, but did not include built-in support for interactive scenarios (i.e, the scenario sidebar, scenario widgets, spotlight scenarios) nor heuristic evaluation. It resembled a more polished version of the first prototype used in subsection 5.1: an editor with a policy-informed model in the sidebar (Fig. 12). We chose this baseline because it mimicked the system design of many existing “copilot” systems that embeds a chatbot in a sidebar alongside a document environment (e.g., Copilot for Microsoft or Gemini in Google Docs) as well as policy authoring setups used in prior literature (Cheong et al., 2024).
@@ 6.1.3. Starting materials
We prepared scenarios, heuristics, and a small amount of starter text for the policy. For each domain, the first author crafted 10 scenarios that represent a realistic conversation between a human user and an AI chatbot, using the same Llama 3.3 instance as PolicyPad to generate the responses. To maximize realism without access to product interaction logs, we sourced topics and language from datasets containing realistic questions in our domains of interest: MENTAT from Lamparth et al. (Lamparth et al., 2025) for mental health, and r/legaladvice-style cases from Cheong et al. (Cheong et al., 2024) for law. We varied the length of scenarios to be 1–5 conversational turns. For multi-turn scenarios, follow-up user messages were written by the first author. For each group of experts, we randomly sampled half the scenarios (5) to assign to the first task, and the rest were assigned to the second. [^12] We note that seeding a policy prototyping session with a different set of scenarios may change its policy outcomes; we discuss how this can be leveraged favourably in this in Section 9. In the system condition, scenarios were loaded directly into the system. In the baseline condition, the first user messages across the 5 scenarios were copied into a Google Doc and shared with participants upon request. [^13] In the baseline condition, we gave participants the option of starting with or without first browsing these user messages.
Besides scenarios, we provided three basic heuristics to encourage clear and precise policy writing that draws from real-world professional practices. We also provided an Objectives section in the policy as examples of policy statements, drawn from objectives in OpenAI’s Model Spec (OpenAI, 2025b). The full starter heuristics and policy are available in Appendix D.
@@ 6.2. Procedure
  • Post-study survey: Group members swapped policies with another group in their domain and rated the policies on 5-point Likert scale questions (see Appendix E).
@@ 6.3.2. Policy novelty analysis (RQ2)
To analyze the novelty of policies—whether they contribute new perspectives, ideas, considerations, dependencies, or approaches to existing policies—we gathered experts’ policy statements from all sessions and evaluated each against publicly available policies for guiding responsible model behavior (the “existing set”). We combined all policies from OpenAI’s Model Spec (OpenAI, 2025b), Claude’s Constitution (Anthropic, 2023), and principles derived from workshops with legal experts in prior work (Cheong et al., 2024), to represent the set of existing policies. [^14] Excluding policies specifically targeted at moderating hate speech and disturbing content (e.g., (Lam et al., 2024)), as they are orthogonal to the policies we focus on in the prototyping sessions.
@@
    For each expert-written policy statement, we used 3 prompts with varying definitions of novelty [^15] Executed on GPT-4.1. Prompts provided in Supplementary Materials. to make binary novelty decisions against the existing policies. Our prompts also required the LLM to generate a justification for its decision.
@@
    Any policies that were not unanimously determined to be novel in all 3 prompt evaluations were considered not novel. For the remaining policies, we further prompted the model to retrieve relevant quotes from the existing policies to be used as context for human evaluation. [^16] Also through GPT-4.1. Retrieved quotes were examined by the first author alongside the full existing set to catch for hallucinations and any missed quotes.
@@
    Finally, two human annotators (members of our research team) reviewed the list of policies and quotes to make a final novelty determination. Annotations were first done independently, with an initial Cohen’s Kappa of 0.41. Disagreements primarily arose from differences in the interpretation of novelty; they were resolved via a round of discussion. If annotators failed to reach a consensus (which was the case for 2 policy statements), the policy statement was considered not novel by default.
While the reliability of LLMs for making content judgments has been called into question by recent work (Thakur et al., 2025; Chehbouni et al., 2025; Szymanski et al., 2025; Kim et al., 2025; Krumdick et al., 2025; Chehbouni et al., 2025), we structured our evaluation process to minimize the potential impacts of these factors through deliberate prompt design. Specifically, we attempted to control for model sensitivity to prompts by using multiple variations of prompts for initial evaluation. Additionally, we incorporated existing policies from prior work (OpenAI, 2025b; Anthropic, 2023; Cheong et al., 2024) into the prompt and request justifications and quote extraction to identify possible model hallucinations. Finally, we ensured the final novelty determination is done by human annotators. Overall, we believe that this process should yield a conservative determination of novelty, while also ameliorating challenges around human attention during review and comparison of exceptionally long texts. Our prompts used in this evaluation are available in our Supplementary Materials.
@@ 7.1.1. Heuristics built common ground and inspired richer policies (system only)
Besides serving as guidelines, heuristics also served as entry points for deeper discussion on key policy topics. The starter heuristic on incorporating real-world professional practices into the policy initiated discussions in MH01, MH02, and MH04 about motivational interviewing (MI), a foundational technique for therapists. Experts then incorporated various aspects of MI into the policy, such as encouraging “summaries of conversation when appropriate” (MH01), and “Repeating or paraphrasing what [the user] is saying” (MH02). MH02 and MH03 brought up limits of confidentiality—when a mental health expert is legally or ethically required to break confidentiality to share client information, even though expert-client conversations are otherwise private. MH02 brainstormed situations in which experts needed to break confidentiality (e.g., “immediate risk of harm to oneself or others; suicidal thoughts, urges, or behaviors; presence or risk of non-suicidal self injury”) and added a policy to “avoid using MI” in those situations. MH03 agreed that models, just like when experts work with clients, should provide a disclaimer early on in the conversation of conditions under which confidentiality will be broken and remind the user that “[the conversation] is not a confidential setting” when those conditions are triggered.
Interestingly, a couple groups of legal experts used the heuristic to debate whether real-world practices for lawyers and other legal professionals should even apply to AI, since they clearly established that AI does not have the same legal status as human lawyers. P17 in L02 shared that while “we lawyers do have rules of professional responsibility that we need to adhere to, they don’t apply to non-lawyers” and suggested removing that heuristic. P20 in L03 echoed that sentiment: “Those ethical rules of lawyers do not apply to AI systems.” Their group member, P19, agreed, and noted that “if ethical rules of lawyers did apply, then the AI model cannot even begin to suggest answers.” In general, this is an important point of distinction between the mental health and legal policies; we unpack this in more detail in our paper’s Discussion.
@@
Participants also expressed appreciation of the ability to edit the response and receive a system-generated policy suggestion. 100% of the policy suggestions were accepted in our studies. P17 shared that the suggestions were helpful in articulating their thoughts: “sometimes it’s difficult to put your thoughts into words, and the [suggestions] are helping you with that.” P10 agreed, saying that “the ability to pull that response in, edit it, and have a generated guardrail could be a huge time saver.” They saw as a way of removing the need for low-level wordsmithing: “We don’t need to edit that response perfectly, but if we can make it clear what our priorities are, and then see if the AI gets our nuance, that’s pretty incredible.” We also observed that in MH04, P9 and P10’s policy drafting began slowly, but accelerated considerably when they received policy suggestions that inspired more ideas. P7 shared why they accepted policy suggestions even when they seemed imperfect: “I liked the policy [suggestions], even if they weren’t necessarily dead on. It gave us ideas for other [policy statements].”